OpenStudy (anonymous):

If f(x) = x - 2, for what value of x does f (3x) = f(x+1)?

Parth (parthkohli):

Okay, so what would f(3x) be?

OpenStudy (anonymous):

you have to solve for it :P u substitute x-2 in for the Y value in the second equation

OpenStudy (anonymous):

then solve for x

Parth (parthkohli):

Hint: f(3x) = 3x - 2 and f(x + 1) = x + 1 - 2

Parth (parthkohli):

You will have f(3x) = f(x + 1) = 3x - 2 = x + 1 - 2 Am I right?

Parth (parthkohli):

Can you solve the following equation then?\[\rm \large3x - 2 = x + 1 - 2 \]

OpenStudy (anonymous):

:o hmmmm why is she asking answer... just tell her the concept :P